OpenStudy (anonymous):

What is the solution to the equation? -18+4+x=9-19

OpenStudy (denonakavro):

x=4

OpenStudy (photon336):

Combine your terms. -18+4 = -14+x = -10 isolate X -14+x = -10 +14 +14 x = 4

OpenStudy (anonymous):

thank you !!